# Interval Regression: A Comparative Study with Proposed Models

### v1

@@ 3.1 Model Configuration
Except for the AFT model in XGBoost, the Squared Hinge Error is used as the evaluation function for both the cross-validation process and model training. All models are implemented using Python, except for the Linear model. The configuration for each model below is applied to a single train/test set pair.
@@
This model predicts a single value from the train set using only target intervals. The value is a constant that minimizes the Squared Hinge Error (1) between this constant and the set of target intervals.
@@
The R package penaltyLearning [20] is used to implement the Max Margin Interval Regression model with L1 regularization. The regularization parameter starts at 0.001, increasing by a factor of 1.2 until no features remain, with cross-validation (cv = 5) on the train set to determine the optimal L1 regularization value.
@@
[…]
@@
The model was implemented using the Python package xgboost. Cross-validation (cv = 5) was performed to select the optimal hyperparameters, following the same grid search described in Barnwal et al. [11]:
@@
The Euclidean distance metric is applied to normalized features (mean 0, standard deviation 1) to determine the nearest neighbors, implemented using the NearestNeighbors module from the sklearn.neighbors package. A cross-validation (cv = 5) is used to select the optimal value of $K$. The candidate values for $K$ range from 1 to $\lceil\sqrt{n}\rceil$, where $n$ is the number of train instances.
MMIF
MMIF is an ensemble of 100 MMITs. For each MMIT, two-thirds of the train dataset is randomly selected for training, while the remaining one-third serves as the out-of-bag (OOB) set for validation. Additionally, one-third of the original features are randomly chosen for each MMIT. A cross-validation (cv = 5) is used to select the optimal hyperparameters for each MMIT including max_depth {2, 5, 10, 15, 20, 25} and min_split_samples {2, 5, 10, 20, 50}. Each MMIT has its own OOB error, which is used to determine the aggregation rule. Let $\mathcal{T}_{i}$ be an MMIT with OOB error $e_{i}$. The weight assigned to $\mathcal{T}_{i}$ is given by $w_{i}=\frac{\frac{1}{e_{i}}}{\sum_{j=1}^{100}\frac{1}{e_{j}}}$ and the final prediction from MMIF is computed as $\sum_{i=1}^{100}w_{i}\mathcal{T}_{i}(\mathbf{x})$ where $\mathbf{x}$ is the set of features.
@@
Figure 3: The mean and standard deviation of test squared hinge errors for datasets with a high number of features (due to significantly higher error, the Constant model is omitted to improve comparative visualization). Tree-based models generally perform exceptionally well due to their inherent feature selection mechanism. On the other hand, while MLP with ReLU activation function is a more generalized Linear model, it fails to outperform the Linear model. One reason for this is that when a dataset contains a majority of noisy features, MLP cannot effectively reduce their impact on predictions in the same way that a Linear model with L1 regularization can.
@@
Table 1: The comparison of the performance and consistency of 7 models across 39 datasets.
4 Discussion
@@
To ensure a fair and comprehensive comparison, the quality of the datasets is crucial. First, the quality of features is considered. The Constant model (or Featureless), being the simplest, is expected to perform the worst across all datasets. However, in 19 datasets, it is not ranked the lowest. Notably, in 8 of these datasets, it is ranked 4th place or higher, suggesting that the features in these datasets provide limited predictive value.
Next, the relationship between features and targets is examined. Since Linear models assume a linear relationship, while MMIT assumes a highly nonlinear one, dataset linearity is assessed based on performance and consistency from both Linear and MMIT models. If the Linear model outperforms MMIT with greater consistency, the dataset is more likely to exhibit a linear relationship. However, datasets where Linear performs better but with lower consistency than MMIT are not considered. From that criteria, out of 39 datasets, 15 are identified as more linear than nonlinear, including autohorse, autompg, bodyfat, breasttumor, cholesterol, cleveland, elusage, fishcatch, fruity, meta, pbc, simulated.linear, sleep, and vineyard. Conversely, 15 datasets are more nonlinear than linear, such as autoprice, basketball, cloud, cpu, lowbwt, lymphoma.mkatayama, machine.cpu, mbagrade, pollution, pwlinear, pyrim, servo, simulated.abs, simulated.sin, and strike. This statistic demonstrates the diversity and quality of the datasets, ensuring a fair comparison across different scenarios.
@@
Let’s examine the performance and consistency rankings in Table 1. Starting with performance, it’s clear that MMIF outperforms other models in most cases. Linear, MMIT and MLP are comparable, while Constant, KNN, and the AFT model in XGBoost show the poorest performance. When it comes to consistency, MMIF stands out as the most consistent. This is expected, as MMIF being an ensemble model, relies on a group of weak MMITs to make the final prediction, which helps improve its stability. Again, Constant, KNN, and the AFT model in XGBoost show the lowest consistency.
@@
As shown in Table 1, there is no universally optimal model for every scenario. However, if complexity is not a primary concern, and the goal is to select the model with the best overall performance and consistency, MMIF is likely the first choice. Alternatively, if model complexity and training time are critical factors, MMIT or Linear are preferable, as they are less complex than MMIF, require significantly fewer training resources, and still maintain reliable performance.
@@
This is a comparison worth considering. MMIF was introduced with the expectation that an ensemble of MMITs would outperform a single MMIT, and the results confirm this hypothesis. Specifically, across 39 datasets, MMIF is found to outperform MMIT in 28 of them, which constitutes the majority. The reason for this is straightforward: MMIF, being an ensemble of MMITs, is able to reduce overfitting by not relying on a single MMIT.
@@
The AFT model in XGBoost is one of the most complex models in this study, but its performance and consistency do not justify the complexity. It has a limitation in that it can only handle non-negative target intervals. To make the model applicable to a general Interval Regression setting, the target intervals must be transformed using exponential function. This introduces a weakness, as the model becomes highly sensitive when the dataset contains many left-censored target intervals. When predictions are made on the exponential scale, if the predicted value is close to zero, the mapping back to the original scale can result in large negative values, as seen in some test cases where predictions like $-10^{6}$ were made.
@@
The experiments conducted in [11] indicate that the AFT model in XGBoost performs among the worst across their six simulated datasets when compared to the Linear model and MMIT.
A potential direction for future work is to transform left-censored intervals into interval-censored intervals by replacing all lower bounds of $\infty$- ∞ with a finite value. This adjustment aims to reduce the model’s sensitivity to left-censored intervals and mitigate the tendency to produce excessively large negative predictions.
@@
[…]

### v2

@@ 3.1 Model Configuration
Except for the AFT model in XGBoost, the Squared Hinge Error is used as the loss function for both the cross-validation process and model training. The configuration for each model below is applied to a single train/test set pair.
@@
This model predicts a single value from the train set using only target intervals. The value is a constant that minimizes the Mean Squared Hinge Error (5) between this constant and the set of target intervals.
@@
The model Max Margin Interval Regression model is implemented with L1 regularization. The regularization parameter starts at 0.001, geometrically increasing by a factor of 1.2 until no features remain, with cross-validation (cv = 5) on the train set to determine the optimal L1 value.
@@
A cross-validation (cv = 5) was used to select the optimal hyperparameters, including max_depth {0, 1, 5, 10, 20, $\infty$} and min_sample {0, 1, 2, 4, 8, 16, 20}.
@@
Cross-validation (cv = 5) was performed to select the optimal hyperparameters, following the same grid search described in Barnwal et al. [14]:
@@
The Euclidean distance metric is applied to normalized features (mean 0, standard deviation 1) to determine the nearest neighbors. A cross-validation (cv = 5) is used to select the optimal value of $K$. The candidate values for $K$ range from 1 to $\lceil\sqrt{n}\rceil$, where $n$ is the number of train instances.
@@
MMIF
MMIF is an ensemble of 100 MMITs. For each MMIT, two-thirds of the train dataset with one-third of the original features are randomly selected for training, while the remaining instances serves as the out-of-bag (OOB) set for validation. A cross-validation (cv = 5) is used to select the optimal hyperparameters for each MMIT including max_depth {2, 5, 10, 15, 20, 25} and min_split_samples {2, 5, 10, 20, 50}. Each MMIT has its own OOB error, which is used to determine the aggregation rule. Let $\mathcal{T}_{i}$ be an MMIT with OOB error $e_{i}$. The weight assigned to $\mathcal{T}_{i}$ is given by $w_{i}=\frac{\frac{1}{e_{i}}}{\sum_{j=1}^{100}\frac{1}{e_{j}}}$ and the final prediction from MMIF is computed as $\hat{y}=\sum_{i=1}^{100}w_{i}\mathcal{T}_{i}(\mathbf{x})$, where $\mathbf{x}$ is the set of features.
@@
Table 4: The comparison of the performance and consistency of 7 models across 39 datasets. For example, the constant model has the worst performance (rank 7th) in 20 out of 39 datasets, model mmif has the best consistency (rank 1st) in 12 out of 39 datasets.
3.2 Results
For each of the 39 datasets, 7 models were implemented, with each model producing 5 hinge error values. Figure 4 shows the comparison on 3 highly noisy simulated datasets, where 95% of the features are noise (19 out of 20). Figure 5 shows the comparison on 6 slightly unstable datasets, where the feature-to-instance ratio is high (from 0.06 to 1.12). Figure 6 shows the comparison on 6 fairly stable datasets, where the feature-to-instance ratio is low (from 0.009 to 0.09). For more detailed comparison of all 39 datasets—including comparisons on hinge loss, confidence intervals, and IQR boxes—see the supplemental material here: SUPPLEMENTAL MATERIALS.
4 Discussions and Conclusions
@@
To ensure a comprehensive comparison, the quality of the datasets is important. The datasets employed have a variety of characteristics, including high- and low-dimensional data, real-world and synthetic sources, linear and nonlinear patterns, as well as low- and high-noise conditions. An overview of the dataset qualities is provided in Table 3.
@@
The performance and consistency rankings are in Table 4, 1st means the best and 7th means the worst. About the performance, it’s clear that MMIF outperforms other models in most cases. Linear, MMIT and MLP are comparable, while Constant, KNN, and the AFT model in XGBoost show the poorest performance. About the consistency, MMIF stands out as the most consistent. This is expected, as MMIF being an ensemble model, relies on a group of MMITs to make the final prediction, which helps improve its stability. Again, Constant, KNN, and the AFT model in XGBoost show the lowest consistency.
@@
As shown in Table 4, there is no universally optimal model for every scenario. However, if complexity is not a primary concern, and the goal is to select the model with the best overall performance and consistency, MMIF is likely the first choice. Alternatively, if model complexity and training time are critical factors, MMIT is preferable, as it’s less complex than MMIF, require significantly fewer training resources, and still maintain reliable performance.
@@
This is a comparison worth considering. MMIF was proposed with the expectation that an ensemble of MMITs would outperform a single MMIT, and the results confirm this hypothesis. Specifically, across 39 datasets, MMIF outperformed MMIT in 28 of them. This makes sense: MMIF, being an ensemble of MMITs, is able to reduce overfitting by not relying on a single MMIT.
@@
The AFT model in XGBoost is the complex models in this study, but its performance and consistency do not justify the complexity. It has a limitation in that it can only handle non-negative target intervals. To make the model applicable to a general Interval Regression setting, the target intervals must be transformed using exponential function. This introduces a weakness, as the model becomes highly sensitive when the dataset contains many left-censored target intervals. When predictions are made on the exponential scale, if the predicted value is close to zero, the mapping back to the original scale can result in large negative values, as seen in some test cases where predictions like $10^{6}$ were made.
@@
The experiments conducted in [14] (Figure 2.b) indicate that the AFT model in XGBoost performs among the worst across their six simulated datasets when compared to the Linear and MMIT.
A potential direction for future work is to transform left-censored intervals into interval-censored intervals by replacing all lower bounds of -$\infty$ with a finite value. This adjustment aims to reduce the model’s sensitivity to left-censored intervals and mitigate the tendency to produce excessively large negative predictions.
@@
Test Error for each model in each dataset
The two models with the highest mean test Hinge Squared Error have been omitted for clearer visualization.
Ranking table of performance and consistency of each model